OpenStudy (anonymous):

Solve by elimination x+3y=-14 2x-4y=32

OpenStudy (alexandervonhumboldt2):

x+3y=-14 2x-4y=32

OpenStudy (alexandervonhumboldt2):

subtract equation 2 from equation 1

OpenStudy (alexandervonhumboldt2):

x+3y-2x+4y=-14-32

OpenStudy (alexandervonhumboldt2):

-x=-46-7y

OpenStudy (alexandervonhumboldt2):

now substitute it in one of equations and solve for y

OpenStudy (alexandervonhumboldt2):

-(-46-7y)+3y=-14

OpenStudy (alexandervonhumboldt2):

46+10y=-14

OpenStudy (alexandervonhumboldt2):

10y=-60 y=-60/10=??????????/

OpenStudy (alexandervonhumboldt2):

after solving for y substitue y into one of the equation and solve for x

OpenStudy (anonymous):

Okay, thanks for you help! :)